Ordered: 5 mcg/kg/minute

Weight: 220 lbs

Available: Standard solution of 1 gram in 250 ml

Find the infusion rate.
Answer #1

ANSWER.

* Calculation of infusion rate(ml/hr).

         ml/hr = Volume(ml) / Time(hr).

Calculation of volume.

       Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

Here

         Ordered dose = 5 mcg/kg/min.

         Weight of the patient = 220 lbs.

Convert lbs to kg for calculation

                   1lbs = 0.454 kg

      Weight of the patient = 220 * 0.454

                                        = 99.9 kg

       Desired dose(D) = 5 * 99.8

                                   = 499 mcg.

       Dose in hand(H) = 1 gm

Convert 1 gm to mcg for calculation

               1 gm = 1000000 mg.

      Dose in hand(H) = 1000000 mcg.

            Quantity(Q) = 250 ml.

As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 499/1000000 * 250

                                                        = 0.0005 * 250

                                                        = 0.13 ml.

                               Time = 1 min.

     Convert ml/min to ml/hr by multiply with 60

                            ml/hr = 0.13 * 60

                                     = 7.8 ml.